Tutorials:

Mini Probability Tutorial Creator of Tutorial: KendrickLamar2014

Tutorials:

Probability is the likelihood that a certain outcome will occur. The probability of an event ranges from 0 to 1. You can write a probability as a fraction, a decimal, or a percent. If an event has a probability of 0, it means it will not happen. If an event has a probability of 1, it means it is certain to happen. For example, the probability of scoring a touchdown in a hockey game is 0. (Touchdowns are not scored in hockey). The probability of snow on the top of Mount Everest is 1. (There is snow on Mount Everest year-round.) When all possible outcomes are equally likely, you can find the theoretical probability of an event using the following formula.\[P(event) =\frac{ number ~of ~favorable ~outcomes }{ number~of~possible~outcomes }\]

Tutorials:

Example: At JFK High School, 400 new students are randomly assigned to an elective class. There are 200 students in art, 100 students in drama, 100 students in PE, and 100 students in graphic design. What is the probability that the student will be assigned to art class? \[P(art) = \frac{ 200 }{ 400 } = \frac{ 1 }{ 2 }\] What is P(graphic design)? \[P(graphic ~design) = \frac{ 100 }{ 400 } = \frac{ 1 }{ 4 }\]